OpenStudy (anonymous):

okay guys I have a calc 2 question: how should I solve this ratio test? Σ k^20 (x)^k /(2k+1)!

OpenStudy (jiteshmeghwal9):

Permutation or combination ?

OpenStudy (nubeer):

is k^20 (x)^k multiplying?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

this is a series but they want to find out the interval and radius of convergence which I know how to figure that out but my main problem is to do the ratio test properly.

OpenStudy (nubeer):

ok for finding the radius of convergence the ratio test should be <1 Lim \[\lim_{k \rightarrow \infty}(ak+1)/ak \] ak =k^20 (x)^k /(2k+1)!

OpenStudy (nubeer):

now to get (ak+1) substitute k by k+1

OpenStudy (anonymous):

okay yeah you get something like \[k^20+1 x^k+1 / (2k+2)\]

OpenStudy (anonymous):

x^k+1 i meant

OpenStudy (nubeer):

hmmm no.. it would be something like this [(k+1)^20 * (x)^k+1] / (2(k+1) + 1)! [(k+1)^20 * (x)^k+1] / (2k+ 3)! do u get it?

OpenStudy (nubeer):

this is ur ak+1 @rayjan

OpenStudy (anonymous):

let me see one min budd

OpenStudy (nubeer):

sure and remember its all this will be written inside mod and its all less then 1 bcos it converges..

OpenStudy (nubeer):

(2k +3 )! can also be written as (2k+3)(2k+2)(2k+1)!

OpenStudy (anonymous):

just a quick question budd about (2k+3)! well the way they explained it to us is that you just add 1 to 2k+1 which would be 2k+2! so I am a little confused there and also how would I cancel my terms I mean what goes away here comparing to original equation?

OpenStudy (nubeer):

no .. it adds 1 into k.. means every k would become k+1

OpenStudy (anonymous):

oh got it so now we have (k+1)^2*x^k+1/(2k+3)! * the original equation so can you please help me to see which terms would cancel?

OpenStudy (nubeer):

yes x^ k+1 can also be written as x^k * x^1..? do u get this?

OpenStudy (anonymous):

ahm no honestly

OpenStudy (nubeer):

i have just break it tell me if u have x^3*x^2 what wil u get?

OpenStudy (anonymous):

x^5

OpenStudy (nubeer):

yes and how u get that? u did x^ (3+2) right?

OpenStudy (anonymous):

right

OpenStudy (nubeer):

so same way.. x^(k+1) can be written x^k * X^1 right?

OpenStudy (anonymous):

yes

OpenStudy (nubeer):

basically when u have same term and its multiplying u just add the powers.

OpenStudy (nubeer):

ok when u apply ratio rule.. x^k will be cancell.. please keep doing on page so u can know what are u doing

OpenStudy (anonymous):

okay one min.

OpenStudy (anonymous):

okay I got it in this form now:\[k^20*x^k/(2k+1)! * (2k+3)!/(k+1)^20*x^k+1\]

OpenStudy (nubeer):